OpenStudy (anonymous):

Which of the following is a solution of x2 + 4x = -15?

OpenStudy (vlery):

what are the following lol

OpenStudy (anonymous):

do you mean \[x ^{2}\]

OpenStudy (anonymous):

x = -2 - sqr 19 x= 2 + sqr 19 x = 2 + isqr 11 x = -2 - 1sqr 11

OpenStudy (anonymous):

Yes it's x^2

OpenStudy (anonymous):

that last option is supposed to say x = -2 - isqr 11

OpenStudy (anonymous):

@Brilliance

OpenStudy (popitree):